OpenStudy (anonymous):

evaluate x2+5(y-3) when x=-3 and y=14

OpenStudy (anonymous):

the only answer close to that is 29, is that what you meant?

OpenStudy (anonymous):

oh hold up i see what i did wrong

OpenStudy (anonymous):

no the choices are a) 64 b)58 c) 61 D)29

OpenStudy (anonymous):

is ur question \[x^2=5(y-3)\]

OpenStudy (anonymous):

x2+5(y-3) x=-3 and y=14 is the question

OpenStudy (anonymous):

-3^2 + 5(14-3) 9+55 64

OpenStudy (anonymous):

okay awesome thanks